OpenStudy (zeesbrat3):

For f of x equals the quotient of the quantity 1 minus x and the quantity 1 plus x and g of x equals the quotient of the quantity x and the quantity 1 minus x, find the simplified form for f [g(x)] and state the domain.

OpenStudy (zeesbrat3):

@Hero @nincompoop @abb0t @kropot72 @Whitemonsterbunny17 @freckles @Miracrown @vera_ewing

OpenStudy (zeesbrat3):

@ganeshie8

OpenStudy (zeesbrat3):

hey

ganeshie8 (ganeshie8):

\[f(x)=\dfrac{1-x}{1+x}\] \[g(x)=\dfrac{x}{1-x}\] like this ?

OpenStudy (zeesbrat3):

Yes

OpenStudy (zeesbrat3):

\[\frac{ 1- \frac{ x }{ 1-x } }{ 1 + \frac{ x }{ 1-x } }\]

ganeshie8 (ganeshie8):

looks good, multply \(1-x\) top and bottom

ganeshie8 (ganeshie8):

\[\frac{ 1- \frac{ x }{ 1-x } }{ 1 + \frac{ x }{ 1-x } }= \frac{ \left(1- \frac{ x }{ 1-x }\right)\left(1-x\right) }{ \left(1 + \frac{ x }{ 1-x }\right)\left(1-x\right) } = \dfrac{1-x-x}{1-x+x}=1-2x\]

OpenStudy (zeesbrat3):

why does it seem so much harder?

ganeshie8 (ganeshie8):

it is not hard if you simply follow the rules

OpenStudy (zeesbrat3):

i see what i did wrong though

OpenStudy (zeesbrat3):

i multiplied the entire fraction off the bottom, not just the denominator

ganeshie8 (ganeshie8):

For domain part, notice that \(g(x)=\dfrac{x}{1-x}\) is undefined at \(x=1\) and since \(1-2x\) is defined for all real numbers, the domain of \(f(g(x))=1-2x\) is all real number except \(1\)

OpenStudy (zeesbrat3):

Thank you

OpenStudy (zeesbrat3):

Can you help with a few more? @ganeshie8

OpenStudy (zeesbrat3):

A particle moves on a line away from its initial position so that after t hours it is s = 2t2 +3t miles from its initial position. Find the average velocity of the particle over the interval [1, 4]. Include units in your answer.